OpenStudy (anonymous):

On a number line, find the midpoint of a segment with the following endpoints: A = -10 and B = 20 1) 10 2) 5 3) 12 4) 5.5

OpenStudy (radar):

-10-----------0----------------------20 midpoint ???? -10----------0-----5---------------20 distance from -10 to 20 on the number is 20-(-10) =30 mid point is 30/2=15 units from either end. -10+15=5 20-15=5 Looks like it is a 5. Choice 2). Is that what you come up with also?